# AuDeRe: Automated Strategy Decision and Realization in Robot Planning and Control via LLMs

### v1

@@ IV-B Baselines
We consider two baseline approaches for comparison with our proposed method. The first baseline is an end-to-end (e2e) approach, referred to as LLM-predict, where the LLM directly predicts the entire control sequence or planning trajectories from the problem description without external computational assistance. The second baseline, denoted as LLM-code, employs the LLM directly generating executable Python code tailored to achieve the specified tasks [7, 6].
@@ IV-C Performance metrics
To systematically evaluate and compare each method (LLM-predict, LLM-code, and our LLM-use-API), we conduct 100 experiments per scenario. Each method is allowed a maximum of six iterative rounds of queries or reprompting in case errors occurred. We collect three key metrics: 1) the success rate, which is the percentage of experiments in which the method successfully completed the task; 2) the average number of iterations to success, representing the average number of query rounds or repromptings needed to achieve a successful outcome; and 3) the iteration-based success rate that shows the success rate as a function of the number of iterations, which provides insights into the efficiency of each method.
@@ IV-E Ablation study
As a further ablation study, we compare our method, which provides API code to the LLM only after the corresponding APIs are called, with an alternative approach where all available API codes are sent directly to the LLM at once. The results shown in Fig. 10 indicate that the two methods perform similarly, with our approach achieving a slightly higher success rate and fewer iterative query rounds.

### v2

@@ IV-B Baselines
We consider two baseline approaches for comparison with our proposed method. The first baseline is an end-to-end (e2e) approach, referred to as LLM-predict, where the LLM directly predicts the entire control sequence or planning trajectories from the problem description without external computational assistance. The second baseline, denoted as LLM-code, employs the LLM directly generating executable Python code tailored to the specified tasks [6, 5]. As discussed in the literature review, these baselines capture the two dominant paradigms of prior LLM-based planning: direct trajectory generation and code generation. Other approaches are typically designed for specific tasks or rely on additional learning frameworks, and thus address different problem formulations outside the scope of our evaluation.
@@ IV-C Performance metrics
To evaluate and compare each method (LLM-predict, LLM-code, and our LLM-use-API), we conduct 100 experiments per scenario. Each method is allowed a maximum of six iterative rounds of queries or reprompting in case errors occurred. We collect three key metrics: 1) the success rate, which is the percentage of experiments in which the method successfully completed the task; 2) the average number of iterations to success, representing the average number of query rounds or repromptings needed to achieve a successful outcome; and 3) the iteration-based success rate that shows the success rate as a function of the number of iterations, which provides insights into the efficiency of each method.
@@ IV-E Ablation study
As a further ablation study, we compare our method, which provides API code to the LLM only after the corresponding APIs are called, with an alternative approach where all available API codes are sent directly to the LLM at once. The results shown in Fig. 10 indicate that the two methods perform similarly, with our approach achieving a slightly higher success rate and fewer iterative query rounds. By exposing only the relevant APIs, our method is more efficient and cost-effective in both API usage and prompt design.